ABC Company's last dividend was $0.6. The dividend growth rate is expected to be constant at 7% for 4 years, after which dividends are expected to grow at a rate of 5% forever. The firm's required return (rs) is 19%. What is its current stock price (i.e. solve for Po)?

Note: Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box. For example, if your answer is $12.345 then enter as 12.35 in the answer box.
Answer #1
Year Dividend
(D*g^n)
PVIF @ 19% PV of Dividneds
1             0.642 0.8403              0.539
2             0.687 0.7062              0.485
3             0.735 0.5934              0.436
4             0.786 0.4987              0.392
Total              1.853

P4 = Price at the end of Year 4

D5 = Dividend for Year 5 = Dividend for Year 4*1.05 = 0.826

Ke = Firm's Required Rate of Return

g = Growth in rate of Dividend

P0 = Current Stock Price

P4 = D5 / (Ke-g) = 0.826 / (0.19-0.05) = 0.826 / 0.14 = 5.90

P0 = PV of Price at the end of Year 4 + PV of Dividends till year 4

= 5.90 / (1.19^4) + 1.853

= 2.942 + 1.853

= 4.795